# Feature-based One-For-All: Heterogeneous Distillation Across Vision Architectures

### v1

@@ Abstract
Knowledge distillation (KD) involves transferring knowledge from a pre-trained heavy teacher model to a lighter student model, thereby reducing the inference cost while maintaining comparable effectiveness. Prior KD techniques typically assume homogeneity between the teacher and student models. However, as technology advances, a wide variety of architectures have emerged, ranging from initial Convolutional Neural Networks (CNNs) to Vision Transformers (ViTs), and Multi-Level Perceptrons (MLPs). Consequently, developing a universal KD framework compatible with any architecture has become an important research topic. In this paper, we introduce a feature-based one-for-all (FOFA) KD framework to enable feature distillation across diverse architecture. Our framework comprises two key components. First, we design prompt tuning blocks that incorporate student feedback, allowing teacher features to adapt to the student model’s learning process. Second, we propose region-aware attention to mitigate the view mismatch problem between heterogeneous architecture. By leveraging these two modules, effective distillation of intermediate features can be achieved across heterogeneous architectures. Extensive experiments on CIFAR, ImageNet, and COCO demonstrate the superiority of the proposed method.
Introduction
Knowledge distillation (KD) has received increasing attention from both academic and industrial researchers in recent years. The goal is to develop a lightweight and efficient student model by transferring knowledge from an already trained, larger teacher model. This knowledge may come in the form of predictions (Zhao et al. 2022; Li et al. 2023), or intermediate features (Chen et al. 2021b, a; Ji, Heo, and Park 2021), from the cumbersome pre-trained teacher model. Despite the significant progress in KD, there remains a gap in addressing cross-architecture distillation comprehensively. Existing KD approaches have predominantly focused on distillation between teacher and student models with homogeneous architectures, particularly feature-based approaches, leaving cross-architecture distillation relatively unexplored. For example, FitNet (Romero et al. 2015) achieves good results in CNN-to-CNN format but performs poorly when distilling in a heterogeneous CNN-to-ViT scenario, only getting 24.06% on CIFAR-100 with ConvNeXt-T teacher and Swin-P student model pair, which is lower than the basic logit-based method KD (Hinton, Vinyals, and Dean 2015) by more than 50%.
@@ Introduction
Figure 1: High-level comparison between former state-of-the-art OFA and our method FOFA. Due to the view mismatch (red lines) and teacher unawareness (brown lines) problems, OFA can only use final logits to supervise student intermediate features (black lines), thereby restricting its utility in downstream tasks. In contrast, FOFA enables feature imitation by solving these two issues with RAA and AFP modules respectively.
@@
Heterogeneous KD, however, offers numerous advantages. First, it helps overcome architecture-specific limitations, allowing the student model to mitigate weaknesses inherent to individual architectures and achieve a better performance than learning from a homogeneous teacher model. For instance, CNNs can assimilate a more comprehensive understanding from ViTs, conversely, ViTs can augment their capacity for local feature extraction by leveraging insights from CNNs. Second, it enhances the robustness, as the student model learns from the varied error patterns and inductive biases of the teacher models, leading to better performance across different tasks and datasets. While a few studies have delved into the realm of heterogeneous architectures (Liu et al. 2022a; Zhao et al. 2023; Ren et al. 2022), it is noteworthy that these architectures often exhibit limitations in accommodating specific teacher-student pairings. As a result, a thorough restructuring of the link between the teacher and student models is essential whenever a new architecture is introduced into the system.
@@
To overcome the existing limitations, OFA-KD (Hao et al. 2023) is the first to propose a generic framework for heterogeneous KD by projecting the intermediate student features into the logits space and subsequently aligning them with the final predictions of the teacher model. The projection into a latent space with reduced architecture-specific details allows the distillation process to concentrate on high-semantic information, thereby enhancing distillation performance across different architectures. However, this projection results in a loss of spatial information that is crucial for downstream tasks. Furthermore, distilling based on the final high-semantic output of the teacher model may not fully exploit the valuable intermediate features.
@@
To effectively leverage the richly informative intermediate layer features and facilitate distillation across diverse architectures, we revisit the fundamental feature-based distillation pipeline to identify the potential issues. As illustrated in Fig. 1, apart from the different perspectives inherent in different architectural biases, which we refer to as the view mismatch problem, teacher unawareness is a notable concern, whereby the teacher model fails to adjust based on the student’s learning progress. Hence, a novel feature-based one-for-all (FOFA) KD framework is proposed for distillation between heterogeneous architectures to tackle these issues. To address the view mismatch problem, a region-aware attention (RAA) module is proposed to let the student model learn how to reblend features from different patches and stages to achieve a similar perspective with the corresponding teacher features via the attention mechanism. Since the perspective is now similar, traditional stage-wise feature alignment across different architectures can be applied normally. For the teacher unawareness problem, an adaptive feedback prompt (AFP) module is introduced. The AFP module employs prompt tuning (Nie et al. 2023) to generate distillation-friendly teacher features with additional branches. Moreover, the difference of the feature map between the student and teacher model, referred to as feedback, is integrated as extra input to the prompt tuning blocks, enabling the teacher model to dynamically adapt its features in response to the student model’s learning trajectory.
@@
  • Extensive experiments manifest that FOFA improves student model performance with gains of up to 16.94% on CIFAR-100 and 3.35% on ImageNet-1K and 3.50% performance gain on COCO.
Related Works
Homogeneous Knowledge Distillation
@@
A majority of the KD methods (Zhao et al. 2022; Park et al. 2019) predominantly concentrate on homogeneous architectures, particularly within feature-based approaches (Romero et al. 2015; Tian, Krishnan, and Isola 2020), as the intermediate features between such architectures exhibit greater similarity, thereby enhancing the efficiency of the distillation process. For instance, ReviewKD (Chen et al. 2021b) introduces a review mechanism that leverages multi-level information from the teacher model to direct the single-level learning of the student model, achieving commendable performance in the distillation process between CNN-based models. The aforementioned methods exhibit efficacy in distillation processes involving homogeneous architectures. Nonetheless, they fail to consider heterogeneous architectures. As such, they may encounter challenges like view mismatch and teacher unawareness when attempting to transfer knowledge from a heterogeneous teacher model.
Generic Heterogeneous Knowledge Distillation
@@
As more model architectures are developed, creating a KD framework that works with any architecture has become crucial. Different architectures have unique biases (Raghu et al. 2021), which require careful design for KD and often limit it to specific teacher-student pairs (Liu et al. 2022a; Zheng et al. 2023; Ren et al. 2022; Abnar, Dehghani, and Zuidema 2020). OFA-KD (Hao et al. 2023) solves this by converting student features into a logits space, allowing alignment across architectures by removing architecture-specific details, but it loses important spatial information in features. In this work, we aim to design a feature-based distillation framework that keeps rich spatial information, improving both classification performance and downstream tasks.
Adaptive Teacher in Knowledge Distillation
@@
In typical KD approaches (Hinton, Vinyals, and Dean 2015; Chen et al. 2021b; Mirzadeh et al. 2020; Li et al. 2023), the teacher model is first trained in isolation and then used to distill the knowledge to the student model. This results in static knowledge transfer that ignores the student’s learning dynamics, often leading to suboptimal outcomes. Drawing inspiration from prompt tuning (Jia et al. 2022), numerous techniques leverage task-specific prompts that can be adjusted to fine-tune the entire teacher model with minimal parameters (Kim, Jang, and Yang 2024; Zhao et al. 2023). For instance, CAKD (Zhao et al. 2023) integrates adaptable prompts into the teacher model, enabling it to manage distillation-specific knowledge while preserving its discriminative capability. In this paper, we propose using prompt tuning in a generic heterogeneous KD framework to create an adaptive teacher. Unlike previous methods that directly feed the teacher feature, we additionally incorporate student feedback to better align the adaptation with the student’s learning process, further improving the performance.
@@
Figure 2: A comprehensive depiction of the general structure of our FOFA framework and proposed modules. (a) RAA: student features across all stages are concatenated and fed into an attention module to learn how to integrate a new feature with a perspective similar to that of the teacher model. (b) AFP: a prompt-tuning method is introduced to modify the output stage features of the teacher model with respect to the student model’s learning process. Note: Only three stages are shown for convenience. In our experiment, all models are split into four stages.
Method
@@
In this section, we first review knowledge distillation (KD) and then introduce our FOFA framework, which addresses existing issues such as view mismatch and teacher unawareness. We describe the adaptive feedback prompt (AFP) and region-aware attention (RAA) components, and summarize the objective function. An overview of the FOFA framework is shown in Fig. 2.
Preliminaries
@@
Knowledge Distillation (KD) (Hinton, Vinyals, and Dean 2015) aims to effectively transfer knowledge from the pre-trained teacher model to a more compact student model. Among the various types of KD, logits and features are the most commonly used forms of knowledge to transfer.
@@
Region-Aware Attention (RAA) Module
@@
T. S. KD DKD DIST OFA FitNet CC RKD CRD FOFA
ArXiv’15 CVPR’22 NerIPS’22 NerIPS’23 ICLR’15 ICCV’19 CVPR’19 ICLR’20
@@
Adaptive Feedback Prompts (AFP) for Teacher
A naive solution is to optimize the teacher model during the distillation process simultaneously (Guo et al. 2020). However, this method requires more computational cost and may lead to catastrophic forgetting (Kemker et al. 2018) in the teacher model. To address this problem, we propose an adaptive feedback prompts (AFP) module, which uses prompt tuning (Jia et al. 2022; Nie et al. 2023) to fine-tune the teacher model’s knowledge with minimal parameters, aiming not only to eliminate information that is not conducive to distillation but also to make adjustments based on student feedback, while preserving the discriminative capacity of the teacher model.
In practice, the AFP modules are inserted before the start of each stage in the teacher model, and each consists of two blocks, prompt block $PB$, and fusion block $FB$, as shown in Fig. 2(b). To accommodate various architectures, we follow (Nie et al. 2023) to develop $PB$ that integrates task-specific knowledge through an additional branch without the need to modify any architecture. Unlike the previous method, which directly feeds the teacher’s features into the $PB$ (Zhao et al. 2023), we introduce $FB$ before the $PB$ to incorporate the difference of the feature map between the student and teacher model, referred to as student feedback, into the teacher model’s features, aiming to make the output prompt able to reflect the student feedback. Specifically, let $M^{AFP}_{i}$ denotes the transformation that transfers the student feature to the same dimension as the teacher’s at the $i$-th stage, we subtract the student features from the teacher features in each stage to create feedback as follows:
$Feedback_{i}=M^{AFP}_{i}(F^{S}_{prev,i})-F^{T}_{prev,i}.$ (5)
$F^{T^{\prime}}_{i}=PB_{i}(FB_{i}(Concat(Feedback_{i},F^{T}_{i})).$ (6)
$L_{Reg}=L_{KL}(p^{T},p^{T^{\prime}}),$ (7)
Overall Objective Function
By integrating the proposed modules into the general knowledge distillation framework, the overall loss function of FOFA is defined as:
$L_{FOFA}=L_{CE}+\alpha L_{KL}+\beta L_{FD}+\gamma L_{Reg},$ (8)
where $L_{CE}$ denotes the standard cross-entropy loss, $L_{KD}$ denotes the KL-divergence loss, and $L_{FD}$ denotes the feature matching loss, $\alpha$, $\beta$, $\gamma$ are trade-off hyperparameters controlling the impact of each loss. For $L_{FD}$, we use hierarchical context loss (HCL) proposed by ReviewKD (Chen et al. 2021b) as a distance function $D$ to provide a more powerful hierarchical context supervision.
Experiments
@@
T. S. KD DKD DIST OFA FitNet CC RKD CRD FOFA
ArXiv’15 CVPR’22 NerIPS’22 NerIPS’23 ICLR’15 ICCV’19 CVPR’19 ICLR’20
@@
Table 3: Results on COCO based on Faster RCNN (Ren et al. 2015) with FPN (Lin et al. 2017). The highest results are indicated in bold, while the second-best results are underlined.
@@
We conduct experiments on various tasks. First, we compare FOFA with other KD methods on two image classification datasets. We then apply our method to object detection to show the pros of feature-based distillation. All results of FOFA are reported by taking the average over 3 trials. Due to the space constraint, please refer to the supplementary for the implementation details.
Datasets
@@
Two image classification datasets and one detection dataset are adopted for our experiments. (1) CIFAR-100 (Krizhevsky, Hinton et al. 2009) is a well-known dataset, containing 50K training and 10K test images of size 32×32 in 100 classes. (2) ImageNet (Deng et al. 2009) is a large-scale classification dataset, which provides 1.2M images for training and 50K images for validation over 1,000 classes. (3) COCO (Lin et al. 2014) is an 80-category object detection dataset, which contains 118K images for training and 5K images for validation.
Models
@@
We evaluate our FOFA on various networks, including ResNet (He et al. 2016), MobileNetv2 (Sandler et al. 2018), and ConvNeXt (Liu et al. 2022b) as CNN-based models, ViT (Dosovitskiy et al. 2021), DeiT (Touvron et al. 2021) and Swin Transformer (Liu et al. 2021) as ViT-based models, and MLP-Mixer (Tolstikhin et al. 2021) and ResMLP (Touvron et al. 2023) as MLP-based models.
Baselines
@@
We compare with various approaches that are feature-based, such as FitNet (Romero et al. 2015), CC (Peng et al. 2019), RKD (Park et al. 2019), and CRD (Tian, Krishnan, and Isola 2020), as well as the baselines that are based on logits, which include KD (Hinton, Vinyals, and Dean 2015), DKD (Zhao et al. 2022), DIST (Huang et al. 2022) and the most recent state-of-the-art (SOTA) OFA-KD (Hao et al. 2023).
Results on Image Classfication

### v3

# Perspective-Aware Teaching: Adapting Knowledge for Heterogeneous Distillation

@@ Abstract
Knowledge distillation (KD) involves transferring knowledge from a pre-trained heavy teacher model to a lighter student model, thereby reducing the inference cost while maintaining comparable effectiveness. Prior KD techniques typically assume homogeneity between the teacher and student models. However, as technology advances, a wide variety of architectures have emerged, ranging from initial Convolutional Neural Networks (CNNs) to Vision Transformers (ViTs), and Multi-Level Perceptrons (MLPs). Consequently, developing a universal KD framework compatible with any architecture has become an important research topic. In this paper, we introduce a perspective-aware teaching (PAT) KD framework to enable feature distillation across diverse architectures. Our framework comprises two key components. First, we design prompt tuning blocks that incorporate student feedback, allowing teacher features to adapt to the student model’s learning process. Second, we propose region-aware attention to mitigate the view mismatch problem between heterogeneous architectures. By leveraging these two modules, effective distillation of intermediate features can be achieved across heterogeneous architectures. Extensive experiments on CIFAR, ImageNet, and COCO demonstrate the superiority of the proposed method. Our code is available at https://github.com/jimmylin0979/PAT.git.
@@
Figure 1: High-level comparison between former state-of-the-art OFA and our method PAT. Due to the view mismatch (red lines) and teacher unawareness (brown dashed lines) problems, OFA can only use final logits to supervise student intermediate features (black solid lines), thereby restricting its utility in downstream tasks. In contrast, PAT enables feature imitation by solving these two issues with RAA and AFP modules respectively.
1 Introduction
Knowledge distillation (KD) has received increasing attention from both academic and industrial researchers in recent years. The goal is to develop a lightweight and efficient student model by transferring knowledge from an already trained, larger teacher model. This knowledge may come in the form of predictions [37, 17], or intermediate features [3, 2, 12], from the cumbersome pre-trained teacher model. Despite the significant progress in KD, there remains a gap in addressing cross-architecture distillation comprehensively. Existing KD approaches have predominantly focused on distillation between teacher and student models with homogeneous architectures, particularly feature-based approaches, leaving cross-architecture distillation relatively unexplored. For example, FitNet [30] achieves good results in CNN-to-CNN format but performs poorly when distilling in a heterogeneous CNN-to-ViT scenario, only getting 24.06% on CIFAR-100 with ConvNeXt-T teacher and Swin-P student model pair, which is lower than the basic logit-based method KD [10] by more than 50%.
@@ 1 Introduction
Heterogeneous KD, however, offers numerous advantages. First, it helps overcome architecture-specific limitations, allowing the student model to mitigate weaknesses inherent to individual architectures and achieve a better performance than learning from a homogeneous teacher model. For instance, CNNs can assimilate a more comprehensive understanding from ViTs, conversely, ViTs can augment their capacity for local feature extraction by leveraging insights from CNNs. Second, it enhances the robustness, as the student model learns from the varied error patterns and inductive biases of the teacher models, leading to better performance across different tasks and datasets. While a few studies have delved into the realm of heterogeneous architectures [20, 38, 29], it is noteworthy that these architectures often exhibit limitations in accommodating specific teacher-student pairings. As a result, a thorough restructuring of the link between the teacher and student models is essential whenever a new architecture is introduced into the system.
@@
To overcome the existing limitations, OFA-KD [8] is the first to propose a generic framework for heterogeneous KD by projecting the intermediate student features into the logits space and subsequently aligning them with the final predictions of the teacher model. The projection into a latent space with reduced architecture-specific details allows the distillation process to concentrate on high-semantic information, thereby enhancing distillation performance across different architectures. However, this projection results in a loss of spatial information that is crucial for downstream tasks. Furthermore, distilling based on the final high-semantic output of the teacher model may not fully exploit the valuable intermediate features.
@@
To effectively leverage the richly informative intermediate layer features and facilitate distillation across diverse architectures, we revisit the fundamental feature-based distillation pipeline to identify the potential issues. As illustrated in Fig. 1, apart from the different perspectives inherent in different architectural biases, which we refer to as the view mismatch problem, teacher unawareness is a notable concern, whereby the teacher model fails to adjust based on the student’s learning progress. Hence, a novel perspective-aware teaching (PAT) KD framework is proposed for distillation between heterogeneous architectures to tackle these issues. To address the view mismatch problem, a region-aware attention (RAA) module is proposed to let the student model learn how to reblend features from different patches and stages to achieve a similar perspective with the corresponding teacher features via the attention mechanism. Since the perspective is now similar, traditional stage-wise feature alignment across different architectures can be applied normally. For the teacher unawareness problem, an adaptive feedback prompt (AFP) module is introduced. The AFP module employs prompt tuning [24] to generate distillation-friendly teacher features with additional branches. Moreover, the difference of the feature map between the student and teacher model, referred to as feedback, is integrated as extra input to the prompt tuning blocks, enabling the teacher model to dynamically adapt its features in response to the student model’s learning trajectory.
@@
  • Extensive experiments manifest that PAT improves student model performance with gains of up to 16.94% on CIFAR-100 and 3.35% on ImageNet-1K and 3.50% performance gain on COCO.
2 Related Works
2.1 Homogeneous Knowledge Distillation
@@
A majority of the KD methods [37, 25] predominantly concentrate on homogeneous architectures, particularly within feature-based approaches [30, 32, 36], as the intermediate features between such architectures exhibit greater similarity, thereby enhancing the efficiency of the distillation process. For instance, ReviewKD [3] introduces a review mechanism that leverages multi-level information from the teacher model to direct the single-level learning of the student model, achieving commendable performance in the distillation process between CNN-based models. The aforementioned methods exhibit efficacy in distillation processes involving homogeneous architectures. Nonetheless, they fail to consider heterogeneous architectures. As such, they may encounter challenges like view mismatch and teacher unawareness when attempting to transfer knowledge from a heterogeneous teacher model.
2.2 Generic Heterogeneous Knowledge Distillation
@@
As more model architectures are developed, creating a KD framework that works with any architecture has become crucial. Different architectures have unique biases [27], which require careful design for KD and often limit it to specific teacher-student pairs [20, 39, 29, 1]. OFA-KD [8] solves this by converting student features into a logits space, allowing alignment across architectures by removing architecture-specific details, but it loses important spatial information in features. In this work, we aim to design a feature-based distillation framework that keeps rich spatial information, improving both classification performance and downstream tasks.
2.3 Adaptive Teacher in Knowledge Distillation
@@
In typical KD approaches [10, 3, 23, 17], the teacher model is first trained in isolation and then used to distill the knowledge to the student model. This results in static knowledge transfer that ignores the student’s learning dynamics, often leading to suboptimal outcomes. This phenomenon has also been validated in [7], where they attempted to improve student model learning by substituting the conventional teacher trained with cross-entropy loss with one trained using MSE loss. Although this approach yields better performance, training an entirely new teacher model often requires significantly higher computational resources. Another research direction draws inspiration from prompt tuning [13], exploring the possibility of leveraging task-specific prompts that can be adjusted to fine-tune the entire teacher model with minimal parameters [15, 38]. For instance, CAKD [38] integrates adaptable prompts into the teacher model, enabling it to manage distillation-specific knowledge while preserving its discriminative capability. In this paper, we propose using prompt tuning in a generic heterogeneous KD framework to create an adaptive teacher. Unlike previous methods that directly feed the teacher feature, we additionally incorporate student feedback to better align the adaptation with the student’s learning process, further improving the performance.
3 Methodology
@@
Figure 2: A comprehensive depiction of the general structure of our PAT framework and proposed modules. (a) RAA: student features across all stages are concatenated and fed into an attention module to learn how to integrate a new feature with a perspective similar to that of the teacher model. (b) AFP: a prompt-tuning method is introduced to modify the output stage features of the teacher model with respect to the student model’s learning process. Note: Only three stages are shown for convenience. In our experiment, all models are split into four stages.
@@
In this section, we first review knowledge distillation (KD) and then introduce our PAT framework, which addresses existing issues such as view mismatch and teacher unawareness. We describe the adaptive feedback prompt (AFP) and region-aware attention (RAA) components, and summarize the objective function. An overview of the PAT framework is shown in Fig. 2.
3.1 Preliminaries
@@
Knowledge Distillation (KD) [10] aims to effectively transfer knowledge from the pre-trained teacher model to a more compact student model. Among the various types of KD, logits and features are the most commonly used forms of knowledge to transfer.
@@
3.2 Region-Aware Attention (RAA) Module
@@
3.3 Adaptive Feedback Prompts (AFP) for Teacher
A naive solution is to optimize the teacher model during the distillation process simultaneously [6]. However, this method requires more computational cost and may lead to catastrophic forgetting [14] in the teacher model. To address this problem, we propose an adaptive feedback prompts (AFP) module, which uses prompt tuning [13, 24] to fine-tune the teacher model’s knowledge with minimal parameters, aiming not only to eliminate information that is not conducive to distillation but also to make adjustments based on student feedback, while preserving the discriminative capacity of the teacher model.
In practice, the AFP modules are inserted before the start of each stage in the teacher model, and each consists of two blocks, prompt block $PB$, and fusion block $FB$, as shown in Fig. 2(b). To accommodate various architectures, we follow [24] to develop $PB$ that integrates task-specific knowledge through an additional branch without the need to modify any architecture. Unlike the previous method, which directly feeds the teacher’s features into the $PB$ [38], we introduce $FB$ before the $PB$ to incorporate the difference of the feature map between the student and teacher model, referred to as student feedback, into the teacher model’s features, aiming to make the output prompt able to reflect the student feedback. Specifically, let $M^{AFP}_{i}$ denotes the transformation that transfers the student feature to the same dimension as the teacher’s at the $i$-th stage, we subtract the student features from the teacher features in each stage to create feedback as follows:
$Feedback_{i}=M^{AFP}_{i}(F^{S}_{prev,i})-F^{T}_{prev,i}.$ (5)
$F^{T^{\prime}}_{i}=PB_{i}(FB_{i}(Concat(Feedback_{i},F^{T}_{i})).$ (6)
$L_{Reg}=L_{KL}(p^{T},p^{T^{\prime}}),$ (7)
3.4 Overall Objective Function
By integrating the proposed modules into the general knowledge distillation framework, the overall loss function of PAT is defined as:
$L_{PAT}=L_{CE}+\alpha L_{KL}+\beta L_{FD}+\gamma L_{Reg},$ (8)
where $L_{CE}$ is the standard cross-entropy loss, $L_{KD}$ denotes the KL-divergence loss between $p^{T}$ and $p^{S}$, and $L_{FD}$ denotes the feature matching loss between $F^{T^{\prime}}$ and $F^{S^{\prime}}$. The weights $\alpha$, $\beta$, and $\gamma$ balance the contribution of each term. For $L_{FD}$, we adopt the hierarchical context loss (HCL) from ReviewKD [3] as the distance function $D$, providing stronger hierarchical supervision.
4 Experiments
[…]
@@ 4 Experiments
T. S. KD DKD DIST OFA FitNet $C$ RKD CRD PAT
@@
T. S. KD DKD DIST OFA FitNet $C$ RKD CRD PAT
@@
Table 3: Results on COCO based on Faster RCNN [28] with FPN [19]. The highest results are indicated in bold, while the second-best results are underlined.
@@
We conduct experiments on various tasks. First, we compare PAT with other KD methods on two image classification datasets. We then apply our method to object detection to show the pros of feature-based distillation. All results of PAT are reported by taking the average over 3 trials. Due to the space constraint, please refer to the supplementary for the implementation details.
@@
Datasets   Two image classification datasets and one detection dataset are adopted for our experiments. (1) CIFAR-100 [16] is a well-known dataset, containing 50K training and 10K test images of size 32×32 in 100 classes. (2) ImageNet [4] is a large-scale classification dataset, which provides 1.2M images for training and 50K images for validation over 1,000 classes. (3) COCO [18] is an 80-category object detection dataset, which contains 118K images for training and 5K images for validation.
@@
Models   We evaluate our PAT on various networks, including ResNet [9], MobileNetv2 [31], and ConvNeXt [22] as CNN-based models, ViT [5], DeiT [34] and Swin Transformer [21] as ViT-based models, and MLP-Mixer [33] and ResMLP [35] as MLP-based models.
@@
Baselines   We compare with various approaches that are feature-based, such as FitNet [30], CC [26], RKD [25], and CRD [32], as well as the baselines that are based on logits, which include KD [10], DKD [37], DIST [11] and the most recent state-of-the-art (SOTA) OFA-KD [8].
4.1 Results on Image Classification